# Debrup Das Work done during internship at Rakuten India Enterprise Private Limited IIT Kharagpur

### v2

@@ 6 Conclusion
We introduce a Tool-augmented Large Language Model (TALM) framework, aka MathSensei, targeted for Mathematical Reasoning. We utilize tools for web-based knowledge retrieval, program execution, and symbolic equation solving. We perform extensive ablations over the individual tools, along with varying the order and combination on complex mathematical reasoning datasets (such as MATH). Our best configuration achieves a $13.5\%$ improvement over GPT-3.5-Turbo (with CoT prompting) on MATH. Our experiments with tool-sequencing methods does not improve over our best configuration. We also observe that benefit of mathematical TALMs are minimal for simpler math word problems (in GSM-8k) and its benefit increases as the required complexity and knowledge for the problem increases through AQuA, MMLU-Math.
@@
We propose a Tool-Augmented LLM framework (TALM), uniquely targeted towards complex mathematical reasoning. Here, we discuss three types of limitations: 1) choice of the set of tools, 2) variants of the PG module for simpler problems and 3) devloping mathematical TALM-specific planning methods.
1.  Here, we choose tools, which intuitively offers knowledge about complex mathematical disciplines and complex equation solving capabilities such as Python with sympy library, Wolfram-Alpha-API and Bing Web Search API. However, we have not explored other solvers which are targeted towards logical complexity or adding commonsense knowledge. In future, a more universal TALM can target adding Z3, SAT solvers and OMCS knowledge base query capabilities.
@@ Acknowledgements
This work is directly supported by Rakuten India Enterprise Private Limited. Additionally, we would like to thank Microsoft Accelerate Foundation Models grant for providing API access to Bing and Azure-OpenAI services.
Appendix A Prompts
Module #FewShot(s)
KR[Uncaptioned image] 7
WA[Uncaptioned image] 5
PG[Uncaptioned image] 7
SG[Uncaptioned image] 4
Table 7: Number of Few-Shot(s) (or In-context example(s)) used in each module.
Figure 4: Overview of the PGRefer to caption Module (without code refinement).
Figure 5: Overview of the Code Refiner module.
Setting Level 1 Level 2 Level 3 Level 4 Level 5
SG ([Uncaptioned image]) 71.8 53.1 41.0 25.6 12.2
WA+SG ([Uncaptioned image]+[Uncaptioned image]) 74.6 60.5 46.6 37.6 21.3
PG+SG ([Uncaptioned image] + [Uncaptioned image]) 83.6 62.4 52.6 40.0 19.8
WA+PG+SG ([Uncaptioned image]+[Uncaptioned image]+[Uncaptioned image]) 76.4 61.5 54.0 40.2 25.2
PG+WA+SG ([Uncaptioned image]+[Uncaptioned image]+[Uncaptioned image]) 79.1 62.8 53.9 41.8 26.9
BS+WA+SG ([Uncaptioned image]+[Uncaptioned image]+[Uncaptioned image]) 74.6 59.3 51.0 35.6 21.0
WA+BS+SG ([Uncaptioned image]+[Uncaptioned image]+[Uncaptioned image]) 76.0 60.1 52.0 39.9 24.6
BS+PG+WA+SG ([Uncaptioned image]+[Uncaptioned image]+[Uncaptioned image]+[Uncaptioned image]) 81.0 60.5 52.9 41.6 25.4
Table 8: Performance of different Settings across varying Levels of Complexity (1-5) on the MATH dataset.
Here, we provide additional information about the prompts used for each LLM-powered module and examples. We first provide the number of few-shot (or in-context) samples used in each module in Table 7. We further provide the example input prompt and output for the Knowledge Retrieval module in Table 9. We provide the example of the Program Generator module without code refinement in Figure 4 and with code refinement in Figure 5.
[Instruction]
Read the following question, generate the background knowledge or relevant mathematical information as the context information that could be helpful for answering the question.
[In-context example(s)]
The smallest distance between the origin and a point on the graph of [$y=\frac{1}{2}x^{2}-9$] can be expressed as $a$. Find $a^{2}$.
Knowledge:
- The question involves the concept of distance between two points in a coordinate system.
- The distance between the origin and a point (x, y) on the graph is given by the formula $\sqrt{(x^{2}+y^{2})}$.
- Minimizing $(x^{2}+y^{2})$ is same as minimizing $\sqrt{(x^{2}+y^{2})}.$
- Substitute $y=1/2x^{2}-9$ into the equation to get $x^{2}+(1/2x^{2}-9)^{2}.$
- Take derivative of above equation and set it to 0.
[Input Question]
What is the following value when expressed as a common fraction: $\frac{1}{2^{1}}+\frac{1}{2^{2}}+\frac{1}{2^{3}}+\cdots+\frac{1}{2^{8}}+\frac{1}{2^{9}}+\frac{1}{2^{10}}$?
@@ Appendix A Prompts
- The question involves understanding of geometric series in mathematics.
- A geometric series is a series of numbers where each term is obtained by multiplying the previous term by a fixed constant, called the common ratio.
- The sum of the first n terms of a geometric series is given by the formula $S_{n}=a(1-r^{n})/(1-r)$, where a is the first term and r is the common ratio.
- In this case, the first term is 1/2 and the common ratio is also 1/2.
- We need to find the sum of the first 10 terms of the series, which can be calculated using the above formula.
Table 9: Example of Knowledge Retrieval (KR[Uncaptioned image]) Prompt   and Output  .
[Instruction]
@@
[Input Question]
If the numbers 4, 5 and 6 are each used exactly once to replace the letters in the expression $A(B-C)$, what is the least possible result?
@@
A.1 Planning Strategies
Plan-And-Solve
Within the Plan-And-Solve (PAS) framework, a dynamic planner (LLM), generates a plan for a given mathematical problem before the start of execution. In our context, the plan consists of the sequence of modules to be run. Notably, this planning approach is inherently non-adaptive, as the strategy lacks the capability to determine the next module based on feedback and the output of the previously executed modules. To instruct the planner LLM, we provide input prompts containing information about each module, along with a few-shot example representing a possible sequence. The prompts utilized for the planner model are detailed in Table 10.
MathSensei with REACT Planner.
The previous modular settings, have a fixed order of execution of the modules. However, we also wish to test out settings where there is power given to the central LLM to call different modules as and when required. This is done by executing (thought, action request, action execution) triplets. The thought serves as a summary of what we have till now in relation to answering the question, the action request is the specific action we wish to take in the next step, and the action execution step calls the necessary module from the modules library to execute the action. We call this setting thought-request-execution (Tho-Re-Exec), an overview of the setup is presented in Figure 2. The results for this setting corresponding to each problem type is presented in Table 11.
Online Resources.
Plan Method Alg P.Cal P.Alg Geom Prob N.Th Int.Alg O.Acc
PAS* 57.3 29.8 65.0 32.4 42.0 47.7 31.9 47.3
REACT* 62.9 30.6 65.1 32.1 42.0 46.1 33.7 48.9
PG + WA + SG ([Uncaptioned image] + [Uncaptioned image] + [Uncaptioned image])* 61.4 32.8 65.2 33.4 45.4 54.2 37.6 50.7
WA + PG + SG ([Uncaptioned image] + [Uncaptioned image] + [Uncaptioned image])* 64.4 32.1 62.8 32.1 46.9 49.4 38.3 50.6
Table 11: Comparison of planning strategies: Plan-And-Solve (PAS) and REACT with two of our best performing settings on 3072 randomly sampled examples from the MATH dataset. Here X* denotes the use of 3072 samples for evaluating method X.
Setting Level 1 Level 2 Level 3 Level 4 Level 5
PAS* 76.0 60.1 53.9 40.5 26.1
REACT* 78.3 62.0 55.4 41.6 27.9
PG + WA + SG ([Uncaptioned image] + [Uncaptioned image] + [Uncaptioned image])* 79.3 65.3 54.3 43.9 31.1
WA + PG + SG ([Uncaptioned image] + [Uncaptioned image] +[Uncaptioned image])* 78.3 65.6 55.8 43.1 30.1
Table 12: Comparing Performance of different Planning Strategies (§LABEL:exp:planning:alldatasets) with two of our Top Performing Settings by varying Difficulty Level of Problems from the MATH dataset.
Subject PG-Exec-Err PG-R-Err SG-Err Tot.Examples
Alg 8 5 2 15
P.Cal 6 9 0 15
P.Alg 4 11 0 15
Geom 3 12 0 15
Prob 8 6 1 15
N.Th 6 7 3 16
Int.Alg 14 0 1 15
Overall Count 51 48 7 106
Table 13: Summary of Error types with PG + SG ([Uncaptioned image] + [Uncaptioned image]) setting on a random subset of 106 examples (MATH dataset); PG-Exec-Err: Code generated by PG[Uncaptioned image] module having syntactical errors; PG-R-Err: Executable python code (from PG[Uncaptioned image]) having reasoning errors; SG-Err: Solution Generator (SG[Uncaptioned image]) alters correct output from PG[Uncaptioned image] to incorrect; Alg: Algebra, P.Cal: Precalculus, P.Alg: Prealgebra, Geom: Geometry, Prob: Probability, N.Th: Number Theory, Int.Alg: Intermediate Algebra. Here we report the absolute count of errors across different subjects.
Module Models/APIs or Tools
KR[Uncaptioned image]
text-davinci-002 ([Uncaptioned image])
text-davinci-003 ([Uncaptioned image])
Llama-2-7B ([Uncaptioned image])
gpt-3.5-turbo ([Uncaptioned image])
BS[Uncaptioned image]
text-davinci-003 ([Uncaptioned image])
gpt-3.5-turbo ([Uncaptioned image])
Bing-Web-Search-API ([Uncaptioned image])
WA[Uncaptioned image]
text-davinci-003 ([Uncaptioned image])
gpt-3.5-turbo ([Uncaptioned image])
Wolfram-Alpha-API ([Uncaptioned image])
PG[Uncaptioned image]
gpt-3.5-turbo ([Uncaptioned image])
Phind-CodeLLama-34B-V2 ([Uncaptioned image])
Python Interpreter ([Uncaptioned image])
CR[Uncaptioned image]
gpt-3.5-turbo ([Uncaptioned image])
SG[Uncaptioned image]
text-davinci-003 ([Uncaptioned image])
gpt-3.5-turbo ([Uncaptioned image])
Table 14: Module Inventory.
Dataset PG-R-Err PG-Exec-Err SG-Err
GSM-8K 18 1 1
AQUA-RAT 13 6 1
Table 15: Case study of error types of PG+SG ([Uncaptioned image]+[Uncaptioned image]) on GSM-8K and AQUA-RAT datasets for 20 randomly selected examples on which PG([Uncaptioned image]+[Uncaptioned image]) fails and SG([Uncaptioned image]) is correct. PG-Exec-Err: Code generated by PG module having syntactical errors; PG-R-Err: Executable python code (from PG ) having reasoning errors; SG-Err: Solution Generator (SG ) alters correct output from PG to incorrect
[Question]
Uriah’s book bag is getting too heavy for him. He needs to remove 15 pounds from it. His comic books weigh 1/4 pound each and his toys weigh 1/2 pound each. If he removes 30 comic books, how many toys does he need to remove?
[PG+SG Output]
Python generator:
from sympy import *
x = symbols(’x’)
eq = Eq((1/4)*(30-x) + (1/2)*x, 15)
print("Equation:",eq)
sol = solve(eq, x)
print("Solution x:",sol)
ans = 30 - sol[0]
print("Toys to remove:",ans)
Python output:
Equation: Eq(0.25*x + 7.5, 15)
Solution x: [30.0000000000000]
Toys to remove: 0
[SG Output]
Solution:
- We know that Uriah needs to remove a total of 15 pounds from his bag.
- He removes 30 comic books, which weigh a total of 30*(1/4) = 7.5 pounds.
- So he still needs to remove 15 - 7.5 = 7.5 pounds.
- Let the number of toys he needs to remove be x.
- The weight of x toys is x*(1/2) = 0.5x pounds.
- We need to solve the equation 0.5x = 7.5 to find x.
- Solving for x, we get x = 15.
- Therefore, Uriah needs to remove 15 toys from his bag.
- the answer is 15
Table 16: Example showing PG+SG([Uncaptioned image]+[Uncaptioned image]) (Incorrect) vs SG ([Uncaptioned image])(correct) for GSM-8K

### v3

@@ 6 Conclusion
We introduce a Tool-augmented Large Language Model (TALM) framework, aka MathSensei, targeted for Mathematical Reasoning. We utilize tools for web-based knowledge retrieval, program generation and execution and symbolic equation solving. We perform extensive ablations over the individual tools, along with varying the order and combination on complex mathematical reasoning datasets (such as MATH). Our best configuration achieves a $13.5\%$ improvement over gpt-3.5-turbo (with CoT prompting) on MATH. Our experiments with tool-sequencing methods does not improve over our best configuration. We also observe that benefit of mathematical TALMs are minimal for simpler math word problems (in GSM-8k) and its benefit increases as the required complexity and knowledge for the problem increases through AQuA, MMLU-Math.
@@
We propose a Tool-Augmented LLM framework (TALM), uniquely targeted towards complex mathematical reasoning. Here, we discuss three types of limitations: 1) choice of the set of tools, 2) variants of the PG module for simpler problems and 3) developing mathematical TALM-specific planning methods.
1.  Here, we choose tools, which intuitively offers knowledge about complex mathematical disciplines and complex equation solving capabilities such as Python with Sympy library, WolframAlpha-API and Bing Web Search API. However, we have not explored other solvers which are targeted towards logical complexity or adding commonsense knowledge. In future, a more universal TALM can target adding Z3, SAT solvers and OMCS knowledge base query capabilities.
@@ Acknowledgements
This work is directly supported by Rakuten India Enterprise Private Limited. Additionally, Debrup Das and PI Somak Aditya are partially supported by the Microsoft Accelerate Foundation Models Research (AFMR) Grant (especially Bing Search API-based experiments have been carried out with the help of AFMR).
Appendix A Planning Experiments
Plan-And-Solve
Within the Plan-And-Solve (PAS) framework, a dynamic planner (LLM), generates a plan for a given mathematical problem before the start of execution. In our context, the plan consists of the sequence of modules to be run. Notably, this planning approach is inherently non-adaptive, as the strategy lacks the capability to determine the next module based on feedback and the output of the previously executed modules. To instruct the planner LLM, we provide input prompts containing information about each module, along with few-shot examples representing a possible sequence. The prompts utilized for the planner model are detailed in Table 22.
MathSensei with REACT Planner.
The previous modular settings, have a fixed order of execution of the modules. However, we also wish to test out settings where there is power given to the central LLM to call different modules as and when required. This is done by executing (thought, action request, action execution) triplets. The thought serves as a summary of what we have till now in relation to answering the question, the action request is the specific action we wish to take in the next step, and the action execution step calls the necessary module from the modules library to execute the action. An overview of the REACT setting applied to the MATH dataset is presented in Fig. 5. The results for this setting corresponding to each problem type is presented in Table 10.
Results.
We evaluate the performance of Plan-And-Solve and REACT on a randomly sampled subset of the MATH dataset of 3100 examples(for which REACT converges). The results show that simple vanilla implementation of the above planners is not sufficient for surpassing our best configuration PG+WA+SG. In particular, the majority of errors for REACT, were as a result of the failure of REACT to converge to a final solution (finish thought state). The variation of the accuracy as a function of the level of the problem (Table 11) shows, REACT* can surpass Plan-And-Solve (PAS) by a small percentage, however it still lags behind our best settings.
{NiceTabular}
p30mm|p35mm Type Error message
Undefined symbols name ’x’ is not defined
Incorrect handling of objects ’FiniteSet’ object has no attribute ’subtract’
Undefined functions name ’divisible_by’ is not defined
Use of libraries without import sympy package not found
Table 12: Syntactic Errors of the PG module.
Unlike planning in traditional closed world setup datasets such as Blocksworld, Logistics, Depot planning, etc., the task of planning in the mathematical reasoning domain presents multiple differences: Firstly, the set of possible actions is not finite as we can query each tool/module with any input string. Moreover, there are no preconditions that need to be satisfied for executing a particular action which makes the planning space much more unbounded. This can lead to long planning chains with (thought,action,execution) triplets where there may be multiple irrelevant actions. As seen from our work, the strengths and limitations of each tool also varies with the type of datasets, subdomains and difficulty levels, which makes the problem non-trivial. Hence, it turns out to be overwhelming to propose a novel planning strategy in this paper. We plan to explore this issue as a future research direction. A planner with a novel architecture and sufficient mathematical knowledge may be required to tackle this aspect.
{NiceTabular}
l|c|ccc|c Dataset Subject PG-Exec-Err PG-R-Err SG-Err Egs.
MATH Alg 8 5 2 15
P.Cal 6 9 0 15
P.Alg 4 11 0 15
Geom 3 12 0 15
Prob 8 6 1 15
N.Th 6 7 3 16
Int.Alg 14 0 1 15
O.Cnt 51 48 7 106
GSM-8K - 1 18 1 20
AQUA - 1 6 13 20
Table 13: Summary of Error types with PG+SG ([Uncaptioned image]+[Uncaptioned image]) setting on a random subset of 106 examples (MATH dataset); for GSM-8K and AQUA we consider 20 random examples, where the setting SG ([Uncaptioned image]) is correct; PG-Exec-Err: Code generated by PG[Uncaptioned image] module having syntactical errors; PG-R-Err: Executable python code (from PG[Uncaptioned image]) having reasoning errors; SG-Err: Solution Generator (SG[Uncaptioned image]) alters correct output from PG[Uncaptioned image] to incorrect; Alg: Algebra, P.Cal: Precalculus, P.Alg: Prealgebra, Geom: Geometry, Prob: Probability, N.Th: Number Theory, Int.Alg: Intermediate Algebra, O.Cnt: Overall Count, Egs.: Examples. Here we report the absolute count of errors across different subjects.
Figure 4: Overview of (a) Python Generator Module and (b) Code Refiner Module
Figure 5: Generated output for example from the MATH dataset for the REACT planning setting.
Figure 6: Distribution of examples where WA+SG (Refer to caption+Refer to caption) is correct and SG (Refer to caption) is wrong, across problem types and level of difficulty (1-5). There are 897 such examples from MATH dataset.
{NiceTabular}
p50mm|p50mm|p50mm Setting: WA+SG ([Uncaptioned image]+[Uncaptioned image])
QUESTION 1: The positive integers up to 2007 are alternately subtracted and added: [12+34++20012002+20032004+20052006+2007.][1-2+3-4$\cdots$+2001-2002+2003-2004+2005-2006+2007.][ 1 - 2 + 3 - 4 + ⋯ + 2001 - 2002 + 2003 - 2004 + 2005 - 2006 + 2007 . ] What is the value of the expression? QUESTION 2: When the expression $-2x^{2}-20x-53$ is written in the form $a(x+d)^{2}+e$, where a, d, and e are constants, what is the sum a+d+e? QUESTION 3: Solve $\frac{2x^{2}+x+3}{x^{2}+x+1}=\frac{2x+1}{x+1}$ for $x$.
[…]
WA API Query:
Sum$[(-1)^{n}*n,\{n,1,2007\}]$
[INCORRECT; since it finds negative of the required answer]
CORRECT Query :
Sum$[(-1)^{(n+1)}*n,\{n,1,2007\}]$ WA API Query:
CompleteSquare[$-2x^{2}-20x-53$]
[CORRECT] WA API Query:
solve $(2x^{2}+x+3)/(x^{2}+x+1)=(2x+1)/(x+1)$ for x
[CORRECT]
WA Output:
Yes, the answer is -1004 .
[Logical errors in LLM generated WA API queries]. WA Output:
Yes, the answer from this object is: $-2(x+5)^{2}-3$
[CORRECT] WA Output:
Yes, the answer is x = -2
[Since WA module returns a single line answer (in many cases), the downstream SG module fails to provide a step-by-step reasoning.]
SG Output:
We use the Wolfram Alpha API to evaluate the expression Sum$[(-1)^{n}*n,{n,1,2007}]$. The answer from Wolfram Alpha is -1004. Therefore, the answer is $\boxed{-1004}$
[INCORRECT] SG Output:
From WolframAlpha response, we find $-2x^{2}-20x-53=-2(x+5)^{2}-3$.Thus we have $a=-2$, $d=-5$, and $e=-3$. So, $a+d+e=-2-5-3=-10$. Therefore, the answer is $\boxed{-10}$
[INCORRECT; wrong interpretation of correct WA response by the downstream SG module. Value of d should be +5 and not -5.] SG Output
We call the WolframAlpha API to solve the given equation.The answer from Wolfram Alpha is x = -2. Therefore, the answer is $\boxed{-2}$
[CORRECT; but lacks reasoning chains]
Table 14: Limitations of the WA Module in WA+SG ([Uncaptioned image]+[Uncaptioned image]) setting.
WA Query
$Sum[i^{k},\{k,1,259\}]$
$Solve[\{725x+727y==1500,729x+731y==1508\},\{x,y\}]$
FactorInteger[3105]
$maximize[y/x,(x-3)^{2}+(y-3)^{2}==6]$
period of $cos(bx)$
Table 15: Examples of generated WolframAlpha (WA[Uncaptioned image]) Queries.
QUESTION 1: Find the remainder when $x^{5}+x^{4}+x^{3}+x^{2}+x$ is divided by $x^{3}-4x$ QUESTION 2: Let $P$ be the point $(0,5)$, let $Q$ be the point $(6,9)$, and let $R$ be the point $(12,0)$. Determine the area of right-angled $\triangle{PQR}$.
[…]
Setting: SG ([Uncaptioned image])
Response:
The result of the division is $x^{2}+5x+\frac{5x}{x^{3}-4x}$
[INCORRECT; Calculation error in intermediate step]
Response:
$QR=\sqrt{(12-6)^{2}+(0-9)^{2}}=\sqrt{180}$
[INCORRECT; Error in distance calculation in intermediate step.]
Setting: WA+SG ([Uncaptioned image]+[Uncaptioned image])
WA API Query:
PolynomialRemainder[$x^{5}+x^{4}+x^{3}+x^{2}+x,x^{3}-4x,x$]
[CORRECT]
WA API Query:
triangle((0,5),(6,9),(12,0))
[CORRECT]
WA Output:
Yes, the answer is $5x^{2}+21x$
[CORRECT]
WA Output:
[…]
[CORRECT]
SG Output:
The answer from Wolfram Alpha is found to be $5x^{2}+21x$. Therefore, the answer is $\boxed{5x^{2}+21x}$
[CORRECT]
SG Output:
[…]
[CORRECT]
Table 16: Utility of the WA module in WA+SG ([Uncaptioned image]+[Uncaptioned image]) setting.
QUESTION: The average of Amys, Ben, and Chris’s ages is 6. Four years ago, Chris was the same age as Amy is now. In four years, Ben’s age will be 35 of Amy’s age at that time. How many years old is Chris now?
[…]
Setting: SG ([Uncaptioned image])
Response:
$\frac{1}{3}(a+b+c)=6$;       $c-4=a$;       $b+4=\frac{3}{5}(a+4)$.
Simplifying the third equation, we get $125b+20=3a+$ 12. Substituting $c=a+4$ into the first equation and simplifying, we get $a+b=10$.
[INCORRECT; Correct equation: 2a +b =14]
Setting: PG+SG ([Uncaptioned image]+[Uncaptioned image])
Response:
from sympy import *
a, b, c = symbols(’a b c’)
eq1 = Eq((a+b+c)/3, 6)
eq2 = Eq(c-4, a)
eq3 = Eq(b+4, Rational(3,5)*(a+4))
sol = solve((eq1, eq2, eq3), (a, b, c))
print("Chris’s age is", sol[c], "years old")
[Leverages the classes Rational, Eq and method solve from Sympy]
PG Output: Chris’s age is 10 years old
[…]
[CORRECT]
Table 17: Utility of the PG module in PG+SG ([Uncaptioned image]+[Uncaptioned image]) setting.
QUESTION: Uriah’s book bag is getting too heavy for him. He needs to remove 15 pounds from it. His comic books weigh 1/4 pound each and his toys weigh 1/2 pound each. If he removes 30 comic books, how many toys does he need to remove?
Setting: SG ([Uncaptioned image])
Response:
We know that Uriah needs to remove a total of 15 pounds from his bag. He removes 30 comic books, which weigh a total of 30*(1/4) = 7.5 pounds. So he still needs to remove 15 - 7.5 = 7.5 pounds.Let the number of toys he needs to remove be x.The weight of x toys is x*(1/2) = 0.5x pounds. We need to solve the equation 0.5x = 7.5 to find x. Solving for x, we get x = 15. Therefore, Uriah needs to remove 15 toys from his bag. The answer is 15. [CORRECT]
Setting: PG+SG ([Uncaptioned image]+[Uncaptioned image])
Response:
from sympy import *
x = symbols(’x’)
eq = Eq((1/4)*(30-x) + (1/2)*x, 15)           [Wrong equation - reasoning error]
print("Equation:",eq)
sol = solve(eq, x)
print("Solution x:",sol)
ans = 30 - sol[0]
print("Toys to remove:",ans)
PG Output: Equation: Eq(0.25*x + 7.5, 15); Solution x: [30.0000000000000]; Toys to remove: 0;
SG Output: From Python code we get that Uriah needs to remove 0 toys. The answer is 0. [INCORRECT]
Table 18: Demonstration of a Reasoning Error committed by the PG module in PG+SG ([Uncaptioned image]+[Uncaptioned image]) setting in a GSM-8K problem.
INSTRUCTION
Read the following question to generate thought, Query for searching Bing Web Search API, that will help to solve the entire problem or specific subproblems in the question.
IN-CONTEXT EXAMPLES
Question: When the expression $-2x^{2}-20x-53$ is written in the form $a(x+d)^{2}+e$, where $a$, $d$, and $e$ are constants, what is the sum $a+d+e$?
Modules used till now: []
Mathematics Problem Type: Algebra
Level of Problem: Level 5
Thought: Since the question involves completing the square let us search how to complete the square.
Query: How do we complete the square of a quadratic equation?
Question: A circle is circumscribed about an equilateral triangle with side lengths of $6$ units each. What is the area of the circle, in square units? Express your answer in terms of $\pi$.
Modules used till now: []
Mathematics Problem Type: Geometry
Level of Problem: Level 5
Thought: To solve this problem, we can search how to find the circumradius of an equilateral triangle.
Query: How to find the circumradius of an equilateral triangle?
[More few shot examples]
INPUT QUESTION
Question: The point $P=(1,2,3)$ is reflected in the $xy$-plane, then its image $Q$ is rotated by $180^{\circ}$ about the $x$-axis to produce $R$, and finally, $R$ is translated by 5 units in the positive-$y$ direction to produce $S$. What are the coordinates of $S$?
Mathematics Problem Type: Precalculus
Level of Problem: Level 3
@@ Results.
Thought: To solve this problem, we can search the web for formulas related to reflection, rotation and translation.
Query: What are the formulas for reflection, rotation, and translation in 3D space?
Table 19: LLM-based Knowledge Retrieval Prompt
INSTRUCTION
Read the following question to generate thought, Query for searching Bing Web Search API, that will help to solve the entire problem or specific subproblems in the question.
IN-CONTEXT EXAMPLES
Question: When the expression $-2x^{2}-20x-53$ is written in the form $a(x+d)^{2}+e$, where $a$, $d$, and $e$ are constants, what is the sum $a+d+e$?
Modules used till now: []
Mathematics Problem Type: Algebra
Level of Problem: Level 5
Thought: Since the question involves completing the square let us search how to complete the square.
Query: How do we complete the square of a quadratic equation?
Question: A circle is circumscribed about an equilateral triangle with side lengths of $6$ units each. What is the area of the circle, in square units? Express your answer in terms of $\pi$.
Modules used till now: []
Mathematics Problem Type: Geometry
Level of Problem: Level 5
Thought: To solve this problem, we can search how to find the circumradius of an equilateral triangle.
Query: How to find the circumradius of an equilateral triangle?
[More few shot examples]
INPUT QUESTION
Question: The point $P=(1,2,3)$ is reflected in the $xy$-plane, then its image $Q$ is rotated by $180^{\circ}$ about the $x$-axis to produce $R$, and finally, $R$ is translated by 5 units in the positive-$y$ direction to produce $S$. What are the coordinates of $S$?
Mathematics Problem Type: Precalculus
Level of Problem: Level 3
OUTPUT
Thought: To solve this problem, we can search the web for formulas related to reflection, rotation and translation.
Query: What are the formulas for reflection, rotation, and translation in 3D space?
Table 20: Bing Web Search Query generator Prompt for Concepts Search
INSTRUCTION
Read the following question to generate thought, answer and final query for searching Wolfram Alpha API, that is being used to solve the problem or specific subproblems.
IN-CONTEXT EXAMPLES
Question: When the expression $-2x^{2}-20x-53$ is written in the form $a(x+d)^{2}+e$, where $a$, $d$, and $e$ are constants, what is the sum $a+d+e$?
Modules used till now: []
Mathematics Problem Type: Algebra
Level of Problem: Level 5
Thought: To find the answer to the problem, we need to rewrite the quadratic expression $-2x^{2}-20x-53$ in the form $a(x+d)^{2}+e$, where a, d, and e are constants. Then, we find the sum a+d+e. Do you know the Wolfram Alpha command to do this?
Answer: Yes, the Wolfram Alpha command to factor the given quadratic expression is: factor$(-2x^{2}-20x-53)$
Final Query: factor$(-2x^{2}-20x-53)$
Question: Harold tosses a nickel four times. What is the probability that he gets at least as many heads as tails?
Modules used till now: []
Mathematics Problem Type: Counting and Probability
Level of Problem: Level 5
Thought: To find the probability that Harold gets at least as many heads as tails when tossing a nickel four times, we can use combinatorial techniques to count the favorable outcomes. However, to simplify the calculations, we can leverage Wolfram Alpha to compute the probability efficiently. Do you know the Wolfram Alpha command to find the probability for this problem?
Answer: Yes, the Wolfram Alpha command to find the probability of getting at least as many heads as tails when tossing a nickel four times is: $Binomial[4,k]*(1/2)^{4}$, where k ranges from 2 to 4 (inclusive).
Final Query: $Sum[Binomial[4,k]*(1/2)^{4},{k,2,4}]$
[More few shot examples]
INPUT QUESTION
Question: Find the constant c such that $(x^{2}-4x+3)(x+5)-(x^{2}+4x-5)(x-c)=0$ for all $x$.
OUTPUT
Thought: To find the constant c such that the given expression is equal to zero for all x, we can expand the expression and collect like terms. Do you know the Wolfram Alpha command to do this?
Answer: Yes, the Wolfram Alpha command to expand the given expression and set it equal to zero is: expand $((x^{2}-4x+3)(x+5)-(x^{2}+4x-5)(x-c))=0$
Final Query: expand $((x^{2}-4x+3)(x+5)-(x^{2}+4x-5)(x-c))=0$
Table 21: Wolfram Alpha API Query generator Prompt
INSTRUCTION
@@
INPUT QUESTION
Question: If the numbers 4, 5 and 6 are each used exactly once to replace the letters in the expression $A(B-C)$, what is the least possible result?